Mozumuhle takes out personal loan of R439200,00 to help finance the building of his holiday house,. The terms of his loan specify equal three monthly repayments over five years, with 13,2%interest per 5,compounded quarterly. The first payment is made three months after the loan was taken out. Find the size of the 3 monthly payment


1
Expert's answer
2021-05-10T18:30:14-0400

We will find it by the formula:

"А=S\times(\frac{r}{(1+r)^n-1})"

S= 439 200

Since the payment is quarterly, we will change the rate and the number of periods accordingly

"r=\frac{0.132}{4}=0.033"

"n=4\times5=20"

"А=S\times(\frac{r}{(1+r)^n-1})=439 200\times(\frac{0.033}{(1+0.033)^{20}-1})=15 852.40"